what is the margin rate currently if I shorted 100 shares of a $50 stock last week, with a 45% margin requirement and today the price is $40
Answer #1

Number of shares shorted - 100

Price of the stock when shorted - 50

Total value of stock - 100 * $50 = 5000

Total margin utilised while shorting the shares 45%. Hence 45% on 5000 = 2,250

Hence in total investment of $5000 your contribution was $2,750 and $2,250 was on margin

Today's value of stock = 100 * $40 = 4000

Now total value of stock would be

you sold at $5000

you can buy as $4000

So difference would be 5000-4000 = 1000

This $1000 would be adjusted in your portfolio. So initial margin was 2250 and the required margin now would be 1250

So Margin rate now would be as below :-

Total Price of the stock - Funds available / Total value of stock = 4000-2750/1250 = 0.3125

= 31.25 %

So the current margin in 31.25 %

Note - Brokerage/Margin fees not added in this calculations